# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 7.4 Operator norm of $R^{(\ell)}_{k,\#}$
As in Section 7.2, the number of important times encoded is to important for our purposes, therefore we improve the way we encode the paths taking into account the tangle-freeness of $\gamma^{\epsilon}_{i}$. We partition the encoding era wise, and partition the important times in each era into three categories: short cycling times, long cycling times and superfluous times. Let $i\in[2m]$ and $\epsilon\in\{\prime,\operatorname{\prime\prime}\}$ be fixed. If it exists, let $(i,t_{0})$ be the smallest time such that $\operatorname{cc}(x_{i,t_{0}+1})\in\{\operatorname{cc}(x_{i,\ast}),\operatorname{cc}(x_{i,\ast+1}),...,\operatorname{cc}(x_{i,t_{0}})\}$ where =1$\ast$=1∗ = 1 if $\epsilon=\prime$ and $\ast=k+1$ otherwise. Let $\ast\leq\sigma\leq t_{0}$ be such that $\operatorname{cc}(x_{i,t_{0}+1})=\operatorname{cc}(x_{i,\sigma})$. By assumption, $\mathrm{C}_{i}=(\operatorname{cc}(x_{i,\sigma}),...,\operatorname{cc}(x_{i,t_{0}+1}))$ will be the only cycle of $G_{\gamma}^{\prime}$ seen by $\gamma_{i}^{\epsilon}$. The last important time $(i,t)\preceq(i,t_{0})$ is called the short cycling time. We denote by $(i,\hat{t})$ the smallest time $(i,\hat{t})\succeq(i,\sigma)$ such that $\operatorname{cc}(x_{i,\hat{t}+1})\notin\mathrm{C}_{i}$. If $\gamma_{i}^{\epsilon}$ stays in $\Gamma_{i,t}$ for its remaining time, by convention we set $\hat{t}=k$ (resp. $\hat{t}=T_{i}+1$) for $\epsilon=\prime$ (resp. for $\epsilon=\prime\prime$). We replace the mark of the short cycling time by $(\bar{y}_{i,t},\vec{x}_{i,t+1},\sigma,\hat{t},\vec{x}_{i,\tau})$ where once again $(i,\tau)$ is such that $\tau\geq t$ and $(i,\tau)$ is the next step outside of $\Gamma_{i,t}$. The time $(i,\tau)$ is the next first, important or merging time after $(i,\hat{t})$. If $\gamma_{i}^{\epsilon}$ stays in $\Gamma_{i,t}$ we set $\tau=k$ (resp. $\tau=T_{i}+1$) for $\epsilon=\prime$ (resp. for $\epsilon=\operatorname{\prime\prime}$). Important times $(i,t^{\prime})$ with $1\leq t^{\prime}\leq t$ or $\tau\leq t^{\prime}\leq k$ (or in the case $\epsilon=\operatorname{\prime\prime}$, $k+1\leq t^{\prime}\leq t$ or $\tau\leq t^{\prime}\leq T_{i}$) are called long cycling times. The other important times with $t<t^{\prime}<\tau$ are called superfluous.
We have our second encoding and we can recover the sequence $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$ from the positions of the merging times, the long cycling times, the short cycling times and their respective marks. For each $i\in[2m]$ and for each $\epsilon\in\{\prime,\operatorname{\prime\prime}\}$, there is at most one short cycling time, one merging time, and as argued in the proof of 7.2 $g^{\prime}+1$ long cycling times. For each time we have $T_{i}\leq\ell$ or $k\leq\ell$ ways to position them. Therefore we have a total of $\ell^{4m(g^{\prime}+3)}$ ways to positions our marked times. We have $|Y_{\gamma}|\leq a+2m=:a^{\prime}$, where the term $2m$ coming from the element $y_{i,k}$ for $i\in[2m]$. We have at most ${a^{\prime}}(s-p)^{2}s^{2}$ ways to mark a long cycling time, at most $a^{\prime}(s-p)^{2}s^{2}\ell^{2}$ ways to mark a short cycling time and ${a^{\prime}}^{2}(s-p)^{2}s^{2}$ ways to mark a merging time. We deduce from that
@@
$\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\leq N^{s-p}cm\ell\Delta_{\#}^{2m(\ell+1)}.$
For $\gamma\in\mathrm{W}_{\ell,\vec{T}}^{\prime}(s,a,p)$ we consider for each $\operatorname{cc}\in\chi_{\gamma}$ the connected and bi-partite graph in the sense 7.6 $\mathcal{G}_{\operatorname{cc}}=(\mathcal{X}_{\operatorname{cc}}\bigsqcup\mathcal{Y}_{\operatorname{cc}},\mathcal{E}_{\operatorname{cc}})$ and its underlying covering tree $\mathcal{T}_{\operatorname{cc}}$ constructed as we did in Section 7.3. We consider $\mathcal{G}$ the union graph as in Section 7.3 and
$\mathcal{F}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{T}_{\operatorname{cc}},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{R}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{R}_{\operatorname{cc}}.$
@@
We give the detail of the proof for the case #=#\#$\emptyset$# = ∅. We start again from (32), valid again here. Using the same notations then in the proof of 7.5, for a fixed choice of $\{(y_{u},x_{v}):\leavevmode\nobreak\ (u,v)\in\mathcal{Y}\times\mathcal{X}\}$, we rewrite and bound the second sum in (32) as follows
@@
[…]
@@
where we applied 7.7 between the first and second line and bounded $M\leq cm\ell$ and $\left|\mathcal{L}^{\prime}(\mathcal{F},L)\right|\leq\binom{L+F-1}{F-1}$. We also used the fact that $M+R+F=2m(\ell+1)$. ∎
@@
We use the terminology used in the proof of 7.11 above. Let $A_{\ast}$ be the set of inconsistent arcs of $A_{\gamma}^{\prime}$ (as defined above in 6.5). As argued in the proof of 7.9,$|A_{\ast}|$ is bounded by four times the number of excess edges plus twice the number of merging times. There are at most $g^{\prime}+2m$ excess edges and $3m$ merging times, hence:
@@
Applying 6.5 we obtain the wanted result. ∎
We can now prove 7.10.
Proof of 7.10.
@@
$S=\sum_{\vec{T},s,a,p}|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}{\max}\left(|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\#}(\gamma^{\prime})\right)\leq N^{\mathfrak{c}}\Delta_{\#}^{2m\ell}.$ (40)
We fix $\ell\geq 0$ and $\vec{T}\in[\ell]^{2m}$ and $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$. We set $g^{\prime}=g^{\prime}(s,a,p)=a-s+p$. Owing to 7.13 and 7.12 we have
$|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq c^{m+g^{\prime}}cm\ell\Delta_{\#}^{2m(\ell+1)}N^{-g^{\prime}}.$
Besides we find $c>0$ such that $\ell$, $a+2m$, $s$ and $(s-p)$ are bounded by $cm\ell$. Then owing to 7.11
@@
$S$ $\leq\sum_{\begin{subarray}{c}s,a,p;\\ g^{\prime}(s,a,p)\geq 0,\leavevmode\nobreak\ p\leq 2g^{\prime}+10\end{subarray}}\Delta_{\#}^{2m(\ell+1)}(cm\ell)^{28m(g^{\prime}+8)+1}N^{-g^{\prime}}c^{m+g^{\prime}}$
$\leq(c\ell m)^{227m}\Delta_{\#}^{2m(\ell+1)}\sum_{g^{\prime}}(c\ell m)^{28mg^{\prime}}N^{-g^{\prime}}$
@@
$S\leq N^{\mathfrak{c}}\Delta_{\#}^{2m\ell}\sum_{g^{\prime}}\left(\frac{L}{N}\right)^{g^{\prime}}$
@@
8 Proof of 5.1
All ingredients are now gathered to prove 5.1. Let $10<\alpha,r<$ 1, $10<c<c^{\prime}<$ 1 and $n\geq 0$ be fixed. We can assume for all that follows that $\ell\leq\frac{\ln(N)}{\ln(d)}$ since otherwise the bound of 6.4 would be larger than $1$. Notice that for $N$ large enough, $c\frac{\ell^{4}(4d)^{3\ell}}{N}\leq c\frac{(4d)^{6\ell}}{N}$. Therefore denoting $\Omega$ the event $(M,Q)$ is $\ell$-tangle free and applying 6.4 and 6.3, for any $C>0$ we have
$\mathbb{P}\left(\|A^{\ell}_{\#|_{\boldsymbol{1}^{\perp}}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\right)$ $\leq\mathbb{P}\left(\|A^{\ell}_{\#|_{\boldsymbol{1}^{\perp}}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\cap\Omega\right)+\frac{(4d)^{6\ell}}{N}$
$\leq\mathbb{P}\left(J\geq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\right)+\frac{(4d)^{6\ell}}{N},$
@@
$J:=\|\underline{A}_{\#}^{(\ell)}\|+\frac{1}{N}\sum_{k=1}^{\ell}\|R_{k,\#}^{(\ell)}\|.$
Owing to 7.1 and 7.10 we find $C=C(n)>0$ such that with probability at least $1-2N^{-c^{\prime}}$
$J$ $\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}+\frac{1}{N}\sum_{k=1}^{\ell}e^{C\ln(N)^{1-\alpha}}\Delta_{\#}^{\ell}\leq e^{C\ln(N)^{1-\alpha}}\rho_{n}^{\ell}\left(1+\frac{\ell}{N}\left(\frac{\Delta_{\#}}{\rho_{n,\#}}\right)^{\ell}\right)$
$\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\#}^{\ell}\left(1+\frac{\ell}{\sqrt{N}}\right)$
where in the second line we used that owing to 4.7 we have $\rho_{n}\geq 1$ and $\rho_{n,r}\geq 1-r$ and we considered $\ell\geq\frac{(1-c^{\prime})\ln(N)}{7K^{\prime}_{\#}\ln(d)}$. Adjusting the constant $C$ and for $N$ large enough we have the wanted result.

### v4

@@ 8.4 Operator norm of $R^{(\ell)}_{k,\mathrm{x}}$
[…]
We have our second encoding and we can recover the sequence $(\bar{\gamma}_{i,t})_{(i,t)\in\boldsymbol{T}}$ from the positions of the merging times, the long cycling times, the short cycling times and their respective marks. For each $i\in[2m]$ and for each $\epsilon\in\{\prime,\operatorname{\prime\prime}\}$, there is at most one short cycling time, one merging time, and as argued in the proof of ˜8.1 $g^{\prime}+1$ long cycling times. For each time we have $T_{i}\leq\ell$ or $k\leq\ell$ ways to position them. Therefore we have a total of $\ell^{4m(g^{\prime}+3)}$ ways to positions our marked times. We have $|Y_{\gamma}|\leq a+2m=:a^{\prime}$, where the term $2m$ coming from the element $y_{i,k}$ for $i\in[2m]$. We have at most ${a^{\prime}}(s-p)^{2}s^{2}$ ways to mark a long cycling time, at most $a^{\prime}(s-p)^{2}s^{2}\ell^{2}$ ways to mark a short cycling time and ${a^{\prime}}^{2}(s-p)^{2}s^{2}$ ways to mark a merging time. We deduce from that
@@
$\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{x}}(\gamma^{\prime})\leq N^{s-p}cm\ell\Delta_{\mathrm{x}}^{2m(\ell+1)}.$
For $\gamma\in\mathrm{W}_{\ell,\vec{T}}^{\prime}(s,a,p)$ we consider for each $\operatorname{cc}\in\chi_{\gamma}$ the connected and bi-partite graph in the sense ˜8.1 $\mathcal{G}_{\operatorname{cc}}=(\mathcal{X}_{\operatorname{cc}}\bigsqcup\mathcal{Y}_{\operatorname{cc}},\mathcal{E}_{\operatorname{cc}})$ and its underlying covering tree $\mathcal{T}_{\operatorname{cc}}$ constructed as we did in Section˜8.3. We consider $\mathcal{G}$ the union graph as in Section˜8.3 and
$\mathcal{F}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{T}_{\operatorname{cc}},\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \mathcal{R}=\bigsqcup_{\operatorname{cc}\in\chi}\mathcal{R}_{\operatorname{cc}}.$
@@
We give the detail of the proof for the case $\mathrm{x}=\mathrm{s}$. We start from
$\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{s}}(\gamma^{\prime})$ $=\sum_{\gamma^{\prime}\sim\gamma}\prod_{\operatorname{cc}\in\chi}\prod_{\begin{subarray}{c}f\in\mathcal{E}_{\operatorname{cc}}\\ f=(u,v)\end{subarray}}(Q^{\ell(f)})_{x^{\prime}_{u},x^{\prime}_{v}}\leq\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}\times\mathcal{X}\end{subarray}}\sum_{(\ell(f))_{f\in\mathcal{E}}\in\mathcal{L}^{\prime}(\mathcal{E},M)}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{E}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$ (48)
Using the same notations than in the proof of ˜8.4, for a fixed choice of $\{(y_{u},x_{v}):\penalty 10000\ (u,v)\in\mathcal{Y}\times\mathcal{X}\}$, we rewrite and bound the second sum in (48) as follows
@@
where the summand $\ast$ is over all $(\ell(f))_{f}\in\mathcal{L}^{\prime}(\mathcal{E},M)$, the summand $\ast\ast$ is over $(\ell(f))_{f}\in\mathcal{L}^{\prime}(\mathcal{F},L)$ and finally the summand $\ast\ast\ast$ is over $(\ell(f))_{f}\in\mathcal{L}^{\prime}(\mathcal{R},M-L)$. For convenience we take $(\kappa+1)^{M-L+R}$ as another upper bound in (38). We have
@@
where we applied ˜8.5 between the first and second line and bounded $M\leq cm\ell$ and $\left|\mathcal{L}^{\prime}(\mathcal{F},L)\right|\leq\binom{L+F-1}{F-1}$. We also used the fact that $M+R+F=2m(\ell+1)$. ∎
@@
[…]
@@
Applying Section˜7.3 we obtain the wanted result. ∎
We can now prove Section˜8.4.
Proof of Section˜8.4.
@@
$S=\sum_{\vec{T},s,a,p}|\mathcal{W}^{\prime}_{\ell,\vec{T}}(s,a,p)|\underset{\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)}{\max}\left(|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{x}}(\gamma^{\prime})\right)\leq N^{\mathfrak{c}}\Delta_{\mathrm{x}}^{2m\ell}.$ (49)
We fix $\ell\geq 0$ and $\vec{T}\in[\ell]^{2m}$ and $\gamma\in\mathrm{W}^{\prime}_{\ell,\vec{T}}(s,a,p)$. We set $g^{\prime}=g^{\prime}(s,a,p)=a-s+p$. Owing to ˜8.10 and ˜8.9 we have
$|\mu^{\prime}(\gamma)|\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})\leq c^{m+g^{\prime}}cm\ell\Delta_{\mathrm{x}}^{2m(\ell+1)}N^{-g^{\prime}}.$
Besides we find $c>0$ such that $\ell$, $a+2m$, $s$ and (sp)(s-p) are bounded by $cm\ell$. Then owing to ˜8.8
@@
$S$ $\leq\sum_{\begin{subarray}{c}s,a,p;\\ g^{\prime}(s,a,p)\geq 0,\penalty 10000\ p\leq 2g^{\prime}+10\end{subarray}}\Delta_{\mathrm{x}}^{2m(\ell+1)}(cm\ell)^{28m(g^{\prime}+8)+1}N^{-g^{\prime}}c^{m+g^{\prime}}$
$\leq(c\ell m)^{227m}\Delta_{\mathrm{x}}^{2m(\ell+1)}\sum_{g^{\prime}}(c\ell m)^{28mg^{\prime}}N^{-g^{\prime}}$
@@
$S\leq N^{\mathfrak{c}}\Delta_{\mathrm{x}}^{2m\ell}\sum_{g^{\prime}}\left(\frac{L}{N}\right)^{g^{\prime}}$
@@
9 Proof of ˜7.1
All ingredients are now gathered to prove ˜7.1. Let $0<\alpha,r<1$, $0<c<c^{\prime}<1$ and $n\geq 0$ be fixed. We can assume for all that follows that $\ell\leq\frac{\ln(N)}{\ln(d)}$ since otherwise the bound of ˜7.2 would be larger than $1$. Notice that for $N$ large enough, $c\frac{\ell^{4}(4d)^{3\ell}}{N}\leq c\frac{(4d)^{6\ell}}{N}$. Therefore denoting $\Omega$ the event $(M,Q)$ is $\ell$-tangle free and applying ˜7.2 and ˜7.1, for any $C>0$ we have
$\mathbb{P}\left(\|A^{\ell}_{\mathrm{x}|_{\boldsymbol{1}^{\perp}}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}\right)$ $\leq\mathbb{P}\left(\|A^{\ell}_{\mathrm{x}|_{\boldsymbol{1}^{\perp}}}\|\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}\cap\Omega\right)+\frac{(4d)^{6\ell}}{N}$
$\leq\mathbb{P}\left(J\geq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}\right)+\frac{(4d)^{6\ell}}{N},$
@@
$J:=\|\underline{A}_{\mathrm{x}}^{(\ell)}\|+\frac{1}{N}\sum_{k=1}^{\ell}\|R_{k,\mathrm{x}}^{(\ell)}\|.$
Owing to Section˜8.1 and Section˜8.4 we find $C=C(n)>0$ such that with probability at least $1-2N^{-c^{\prime}}$
$J$ $\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}+\frac{1}{N}\sum_{k=1}^{\ell}e^{C\ln(N)^{1-\alpha}}\Delta_{\mathrm{x}}^{\ell}\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}\left(1+\frac{\ell}{N}\left(\frac{\Delta_{\mathrm{x}}}{\rho_{n,\mathrm{x}}}\right)^{\ell}\right)$ (50)
$\leq e^{C\ln(N)^{1-\alpha}}\rho_{n,\mathrm{x}}^{\ell}\left(1+\frac{\ell}{\sqrt{N}}\right)$
where in the second line we used that owing to ˜6.4 we have $\rho_{n,\mathrm{s}}\geq 1$ and $\rho_{n,\mathrm{c}}\geq 1-r$ and we considered $\ell\geq\frac{(1-c^{\prime})\ln(N)}{7K_{\mathrm{x}}\ln(d)}$. Adjusting the constant $C$ and for $N$ large enough we have the wanted result.
Acknowledgments